how can i slope m and one point on the graph given this equation m=-5/4;(-1,3)
Answer by algebrahouse.com(1659)

m = -5/4 , point = (-1,3)

Point-slope form for the equation of a line is y - y1 = m(x - x1
m = slope
(x1 , y1) are the coordinates of the point

y - 3 = (-5/4)[x - (-1)] {substituted into point-slope form}
y - 3 = (-5/4)(x + 1) {simplified inside parentheses}
y - 3 = (-5/4)x - (5/4) {used distributive property}
y = (-5/4)x + (7/4) {added 3 to each side}

For more help from me, visit: www.algebrahouse.com